### v2

# Throwing Vines at the Wall: Structure Learning via Random Search

Abstract
Vine copulas offer flexible multivariate dependence modeling and have become widely used in machine learning. Yet, structure learning remains a key challenge. Early heuristics, such as Dissmann’s greedy algorithm, are still considered the gold standard but are often suboptimal. We propose random search algorithms and a statistical framework based on model confidence sets, to improve structure selection, provide theoretical guarantees on selection probabilities, and serve as a foundation for ensembling. Empirical results on real-world data sets show that our methods consistently outperform state-of-the-art approaches.
1 Introduction
Copulas provide a flexible framework for modeling complex multivariate distributions by disentangling marginal behavior from the dependence structure (see, e.g., nelsen2007introduction; joe2014). Among them, the class of vine copulas (bedfordcooke2001; aasczadofrigessibakken2009) has become particularly popular in machine learning because it balances flexibility with tractability better than more classical copula families. Recent applications span a wide range of areas, including domain adaptation (Lopez-Paz2013), variational inference (Tran2015), causal inference (Lopez-Paz2016), clustering (tekumalla2017vine), Bayesian optimization (botied), conformal prediction (park2025semiparametric), and generative modeling (kulkarni2018nonparametric; Konstantelos2018; sun2019; Tagasovska2018b).
A vine copula on $d$ variables consists of two components: its vine structure, a nested sequence of undirected trees, and d(d1)/2d(d-1)/2 bivariate (conditional) copulas, called pair-copulas, indexed by its edges. Vine copulas are most popular on problems from a handful to a few dozen variables. In this regime, the number of pair-copulas remains tractable, while the flexibility of the model can capture complex dependence patterns. However, there exists $2^{(d-3)(d-2)/2-1}d!$ possible vines on $d$ variables (morales2011count; joe2011regular), which is of the same order as the number of directed acyclic graphs on $d$ nodes. This super-exponential growth makes exhaustive search infeasible beyond a few variables. As a result, nearly all applications rely on greedy heuristics such as the spanning tree algorithm of Dissmann2013 or the forward selection procedure of kraus2017d.
Despite their simplicity and lack of theoretical foundation, numerous attempts at improving on them had very limited success (see Section˜2.4). The recent review of Czado2022 concludes that such heuristics still represent the state of the art, and that effective structure selection remains an important open challenge in vine methodology.
Figure 1: Average improvement in out-of-sample negative log-likelihood over Dissmann2013.
Contributions
This paper challenges the prevailing view that the standard heuristics are difficult to improve upon. In particular:
  • We propose a vine structure learning algorithm based on hold-out random search and show that it outperforms state-of-the-art methods across a range of real-data settings for tasks ranging from generative modeling to regression.
  • We integrate random search with a state-of-the-art algorithm for constructing model confidence sets (kim2025locally) tailored to vine structures, prove its validity in our context, and provide an efficient implementation. This enables in-sample assessment of whether alternatives outperform a baseline heuristic and yields a principled set of competitive models.
  • We demonstrate that vine regression methods based on averaging over the model confidence set consistently outperform single-vine approaches, including the current state of the art.
In Figure˜1, we illustrate the potential for improvement by comparing the out-of-sample negative log-likelihood of Dissmann2013 to that of vines sampled uniformly at random, selecting the best candidates using an independent validation set (see Section˜4 for details). From the picture, it is clear that the greedy heuristic is suboptimal, and that random search performance improves with the number of candidates.
The new algorithms are conceptually simple, straightforward to implement, and provide immediate benefits for applications of vine copulas in machine learning and related fields.
Outline
In Section˜2, we review the necessary background on vine copulas and structure learning. In Section˜3, we introduce our new structure learning algorithms. In Section˜4, we present an extensive empirical evaluation of our methods and compare them to state-of-the-art approaches. Finally, we conclude in Section˜5.
2 Background
2.1 Copulas
Mathematically, a copula is a multivariate distribution with standard uniform margins. From sklar1959, a random vector $\bm{X}\in\mathds{R}^{d}$ has joint distribution $F$ with univariate marginal distributions $F_{1},\dots,F_{d}$, if and only if there exists a copula $C$ such that, for each $x\in\mathds{R}^{d}$,
$F\left(x_{1},\,\ldots,\,x_{d}\right)=C\left\{F_{1}(x_{1}),\,\ldots,\,F_{d}(x_{d})\right\},$
which is unique if the margins are continuous. Taking partial derivatives on both sides of this equation, the joint density is the product of marginal densities $f_{j}=\partial F_{j}/\partial x_{j}$ and the copula density $c=\partial^{d}C/\partial u_{1}\cdots\partial u_{d}$:
$f(x_{1},\dots,x_{d})=c\left(F_{1}(x_{1}),\dots,F_{d}(x_{d})\right)\times\prod_{j=1}^{d}f_{j}(x_{j}).$
Taking logarithm on both sides further implies that the joint log-likelihood is the sum of the marginal and copula log-likelihoods, which can be exploited in a two-step estimation procedure (genest1995semiparametric; joe1996estimation): first estimate each of the marginal distributions, and then the copula based on the pseudo-observations $\widehat{U}_{i}=\widehat{F}_{i}(X_{i})$, $i=1,\dots,d$.
2.2 Vines
Following the seminal work of joe1996 and bedfordcooke2001; Bedford2002, any copula density $c$ can be decomposed into a product of d(d1)/2d(d-1)/2 bivariate (conditional) copula densities. The order of conditioning in this decomposition can be organized using the following graphical structure.
12341,22,33,41,3||22,4||31,22,33,41,3||22,4||31,4||2,3
Figure 2: A vine structure on four variables.
Definition 2.1.
A vine is a sequence of spanning trees $(V_{t},E_{t})$ with $V_{t}$ and $E_{t}$ respectively the nodes and edges for $t=1,\dots,d-1$, satisfying the following conditions:
  1. (i)
    $V_{1}=\{1,\dots,d\}$.
  2. (ii)
    $V_{t}=E_{t-1}$ for $t\geq 2$.
  3. (iii)
    (Proximity condition) If two nodes in $(V_{t+1},E_{t+1})$ are joined by an edge, the corresponding edges in $(V_{t},E_{t})$ must share a common node.
A simple example of a vine tree sequence is shown in Figure˜2. A vine copula identifies each edge $e$ of the vine structure, with a label $\{j_{e},k_{e}|D_{e}\}$ and a bivariate copula $C_{j_{e},k_{e}|D_{e}}$, called a pair-copula. The sets $\{j_{e},k_{e}\}\subset\{1,\ldots,d\}$ and $D_{e}\subset\{1,\ldots,d\}\setminus\{j_{e},k_{e}\}$ are called the conditioned set and conditioning set, respectively, and we refer to Czado2022 for a precise definition. As for $C_{j_{e},k_{e}|D_{e}}$, it describes the dependence between $U_{j_{e}}$ and $U_{k_{e}}$ conditionally on the subvector $\bm{U}_{D_{e}}$ corresponding to the conditioning set $D_{e}$. The corresponding copula density for $\bm{u}\in[0,1]^{d}$ then factorizes as
$c(\bm{u})$ $=\prod_{t=1}^{d-1}\prod_{e\in E_{t}}c_{j_{e},k_{e}|D_{e}}\bigl(u_{j_{e}|D_{e}},u_{k_{e}|D_{e}}\mid\bm{u}_{D_{e}}\bigr),$ (1)
with $u_{j_{e}|D_{e}}=C_{j_{e}|D_{e}}(u_{j_{e}}|\bm{u}_{D_{e}})$, $C_{j_{e}|D_{e}}$ the conditional distribution of $U_{j_{e}}$ given $\bm{U}_{D_{e}}$, and similarly for $u_{k_{e}|D_{e}}$. As an example, decomposing the Gaussian copula density as in (1) leads to Gaussian pair-copulas $c_{j_{e},k_{e}|D_{e}}$ with parameter equal to the partial correlation coefficient $\rho_{j_{e},k_{e};D_{e}}$.
Thanks to the proximity condition, $u_{j_{e}|D_{e}}$ for any edge $e\in E_{t+1}$ can be computed recursively from the pair-copulas in trees $1,\dots,t$. This way, the parameters of vine copulas can be estimated by iterating between parameter estimation and conditional distribution evaluations, going from the first tree to the last (aasczadofrigessibakken2009). A similar decomposition holds when some variables are discrete, effectively by replacing derivatives of the distribution function by finite differences (panagiotelis2012; funk2025towards).
2.3 Why the Structure Matters
Equation (1) provides a general decomposition of any copula density, regardless of the vine structure. The key to the practical success of vine copulas is a simplifying assumption that the conditional copulas in (1) do not depend on the conditioning value $\bm{u}_{D_{e}}$. This reduces the model to a collection of bivariate copulas, which are straightforward to estimate, yet still highly flexible. Each pair-copula can be chosen independently, allowing complex dependence patterns with varying strength, asymmetry, and non-linear tail behavior. More in-depth discussions of the assumption can be found in stoeber2013simplified; spanhel2019simplified; nagler2025simplified.
The simplifying assumption also elevates the importance of the vine structure itself: different structures can yield substantially different approximations of the true copula density. Finding a structure that fits the data well is therefore crucial, both for predictive performance and for interpretability.
2.4 Related Work
Several previous works have addressed structure learning in vine copula models.
General-purpose methods
The standard approach was introduced by Dissmann2013, whose greedy algorithm constructs a maximum spanning tree based on the absolute value of Kendall’s $\tau$. Variants using information criteria or goodness-of-fit $p$-values (czado2013selection) offer no clear benefits but incur higher computational cost. Attempts to incorporate tests of the simplifying assumption (Kraus2017) also failed to consistently improve performance. More elaborate search strategies based on MCMC (gruber2015sequential; gruber2018) or neural networks (sun2019) can yield gains but are prohibitively expensive: even in the low-dimensional examples provided in the papers, the algorithm run times are on the order, or larger than, those of exhaustive search. Randomized search variants have also been explored. Sampling from common variable orders (zhu2020common) does not improve out-of-sample performance on real data and the Monte Carlo tree search approach with lookahead of chang19a only applies to truncated Gaussian models. Overall, the greedy heuristic of Dissmann2013 remains the default in practice (Czado2022).
Regression problems
Specialized algorithms have been developed for regression tasks. These methods restrict attention to structures that admit closed-form conditional distributions, often star- or path-shaped trees (D- and C-vines). Some extend Dissmann’s approach (chang2019prediction; chang19a; cooke2019vine), while others use forward variable selection with early stopping (kraus2017d; Schallhorn2017; zhu2021simplified; tepegjozova2022nonparametric). Their applicability is thus narrower than general-purpose methods.
Sparse models
Another line of work focuses on sparsity, i.e., identifying independence copulas among pairs, which becomes necessary in higher dimensions to keep models tractable. Approaches include selecting truncation or thresholding levels (kurowicka2011optimal; BrechmannCzadoAas2012; brechmann2015truncation; joe2018parsimonious; nagler2019model), typically combined with spanning-tree heuristics. A different perspective exploits connections between vine copulas and Gaussian DAGs (mueller2018representing; muller2019dependence; muller2019selection), aiming primarily to detect conditional independencies rather than fully optimizing the vine structure. These questions are important but somewhat orthogonal to the focus of this paper, which is the structure itself, rather than the pair-copulas.
3 Methodology
3.1 Setup
Let $\mathcal{D}=\{\bm{Z}_{i}:1\leq i\leq n\}$ be a dataset of $n$ i.i.d. samples in $d$ dimensions, with $\bm{Z}_{i}=(Z_{i,1},\ldots,Z_{i,d})\in\mathds{R}^{d}$. In generative modeling tasks, these are the features of interest, while in regression-like problems we consider $\bm{Z}=(\bm{X},Y)$ as a feature–label pair.
Given a vine structure $\mathcal{V}$ and a dataset $\mathcal{D}$, we assume access to an algorithm that fits an estimated joint density $\widehat{f}_{\mathcal{V},\mathcal{D}}\in\mathcal{F}$, with $\mathcal{F}$ the space of probability densities in $\mathds{R}^{d}$. This density is composed of estimated marginal distributions and a simplified vine copula with structure $\mathcal{V}$.
To evaluate different models, let $L\colon\mathcal{F}\times\mathds{R}^{d}\to\mathds{R}$ denote a loss functional, where $L(f,\bm{Z})$ quantifies the quality of $f\in\mathcal{F}$ based on an observation $\bm{Z}$. Common choices include the negative log-likelihood $L(f,\bm{Z})=-\log f(\bm{Z})$ for generative modeling, or similarly $L(f,(\bm{X},Y))=-\log f(Y|\bm{X})$ for regression-type problems, and we refer to gneiting2014probabilistic for a comprehensive review in the context of probabilistic forecasting.
The expected loss is $L(f)=\mathds{E}_{\bm{Z}}[L(f,\bm{Z})]$, and the goal of model selection is to find, among a set of candidates, the one with minimal expected loss. In practice however, $L(f)$ is unknown and must be estimated from data, which can be done, e.g., via hold-out validation or cross-validation (stone1974cross).
3.2 Random Search Algorithm
Algorithm 1 Hold-out Random Search for Vines.
1:Data set $\mathcal{D}$, hold-out ratio $\eta$, number of candidate models $M$.
2:Optimal vine structure $\widehat{\mathcal{V}}$.
3:Split the data set into disjoint training and validation parts $\mathcal{D}=\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{val}}$ with $\lfloor\eta n\rfloor$ validation observations.
4:Generate candidate structures $\Theta=\{\mathcal{V}_{1},\dots,\mathcal{V}_{M}\}$.
5:For each candidate $\mathcal{V}\in\Theta$:
6:   Use the training set to fit a density $\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}}$.
7:   Compute the validation loss
$L_{\text{val}}(\mathcal{V})=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{\bm{z}\in\mathcal{D}_{\text{val}}}L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{z}).$
8:Select the structure with minimal validation loss:
$\widehat{\mathcal{V}}=\operatorname*{arg\,min}_{\mathcal{V}\in\Theta}L_{\text{val}}(\mathcal{V}).$
To identify a suitable vine structure, we propose a simple and easy-to-implement algorithm based on hold-out random search, summarized in Algorithm˜1. The procedure consists of splitting the data into training and validation sets, generating candidate structures at random, fitting them on the training data, and selecting the one with the smallest validation loss.
For a typical vine model $\widehat{f}_{\mathcal{V},\mathcal{D}}$, both training and inference incurs an $O(d^{2})$ cost per observation, where the factor $d^{2}$ comes from the quadratic number of pair-copulas in the model. This leads to an overall runtime complexity of $O(Mnd^{2})$ for Algorithm˜1. Note that it is embarrassingly parallel over the $M$ candidates, as well as, for a single candidate, tree-wise over pair-copulas. Furthermore, one can reduce the dependence on $d$ from quadratic to linear by truncating the vine at a fixed level (e.g., BrechmannCzadoAas2012).
For generating vine structures, we propose to sample uniformly at random using the algorithm of joe2011regular; see also joe2014 and its implementation in the software libraries VineCopula and pyvinecopulib (vinecopula2025; pyvinecopulib). Other sampling schemes, potentially involving the training data, could be considered as well and also fit the theoretical framework and results ahead; see Section˜D.3 for a sampling scheme based on local perturbations of the Dissmann2013 structure.
While the approach naturally extends to cross-validation, we argue that the additional computational budget is better invested in evaluating more random candidates; see Figure˜1 and Section˜4. In addition, cross-validation does not solve the problem that the selected model may be statistically indistinguishable from other candidates, which we address in the next subsection.
3.3 Vine Confidence Sets
On any given data set, we may want to assess whether the model found by Algorithm˜1 is better than the benchmark heuristic of (Dissmann2013). When this is not the case, we might prefer the benchmark model for qualitative reasons such as ease of interpretability or easier comparison to previous analyses. And even if the benchmark is outperformed, multiple candidates from the random search may be statistically indistinguishable. We tackle this problem via model confidence sets (MCS) (hansen2011model); roughly speaking, subsets of the candidates that contain the “best ones” with high probability.
Define the set of optimal models according to expected out-of-sample loss as
$\Theta^{*}=\operatorname*{arg\,min}_{\mathcal{V}\in\Theta}\mathds{E}[L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})\mid\mathcal{D}_{\text{train}}],$
explicitly allowing for the possibility of multiple ones that are equally good. We emphasize that, in our setup, the goal is not to identify a fixed ground truth. We aim to find structures that lead to trained models with minimal out-of-sample loss. In this setting, the optimal set $\Theta^{*}$ is itself random and determined by the training data, estimation algorithm, and candidate set. In particular, we do not assume that the data was generated from a vine copula model. An MCS is then another random subset that contains each of the optimal models with high probability, as formalized in the following definition.
Definition 3.1 (MCS for Vines).
An $\alpha$-level MCS is a random set $\widehat{\Theta}\subseteq\Theta$ such that
$\forall\,\mathcal{V}\in\Theta^{*}\colon\;\mathbb{P}(\mathcal{V}\in\widehat{\Theta}\mid\mathcal{D}_{\text{train}})\geq 1-\alpha.$
Such an MCS offers marginal guarantees, in the sense that the probabilistic statement applies to each optimal model separately, rather than to the set as a whole. An alternative concept is a uniform MCS requiring that $\Theta^{*}\subseteq\widehat{\Theta}$ with high probability. While this is a stronger guarantee, it is more difficult to achieve in practice, and may lead to unnecessarily large sets including many suboptimal models.
3.4 A Vine MCS Algorithm
Algorithm 2 MCS for Vines.
1:Data set $\mathcal{D}$, hold-out ratio $\eta$, number of candidate models $M$, confidence parameter $\alpha$.
2:Model confidence set of vine structure $\widehat{\Theta}$.
3:Follow Steps 1-3 from Algorithm˜1.
4:Compute losses for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$
$L_{i,m}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z}_{i}),\quad\bm{Z}_{i}\in\mathcal{D}_{\text{val}}.$
[…]
6:Return the vine MCS $\widehat{\Theta}=\{\mathcal{V}_{m}\colon m\in\widehat{\mathcal{I}}\}$.
Several methods for estimating MCSs have been proposed in the literature (hansen2011model; lei2020cross; mogstad2024inference; zhang2024winners; kissel2022black). They generally rely on discrete argmin inference, where the null hypothesis for the optimality of a given candidate $\mathcal{V}\in\Theta$ can be written as
$H_{0,\mathcal{V}}\colon\;\mathds{E}[L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})]\leq\min_{\mathcal{W}\in\Theta}\mathds{E}[L(\widehat{f}_{\mathcal{W},\mathcal{D}_{\text{train}}},\bm{Z})].$
If $\widehat{T}_{\mathcal{V}}$ is a statistic satisfying $\widehat{T}_{\mathcal{V}}\to N(0,1)$ under $H_{0,\mathcal{V}}$, a natural MCS estimator obtains by inverting the test:
$\widehat{\Theta}=\{\mathcal{V}:\widehat{T}_{\mathcal{V}}\leq\Phi^{-1}(1-\alpha)\,\},$
where $\Phi$ is the standard normal distribution function. In this work, we use the DA-test of kim2025locally because of its computational simplicity and good empirical performance. A full description of the test can be found in Appendix˜A for completeness, along with a runtime comparison with alternative methods.
Algorithm˜2 describes how this MCS method fits into our vine structure learning pipeline. The algorithm has runtime complexity $O(Mnd^{2}+Mn)$, where the first term is from the first three steps (see the discussion in Section˜3.2), the second from the MCS procedure, and we refer to Section˜A.2 for this part of the complexity analysis. It also comes with the theoretical guarantee of Proposition˜3.2, whose proof is given in Appendix˜B.
Proposition 3.2.
Suppose
$\limsup_{n\to\infty}\max_{\mathcal{V}\in\Theta}\mathds{E}[|L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}\mid\mathcal{D}_{\text{train}}]<\infty.$
Then Algorithm˜2 yields
$\liminf_{n\to\infty}\min_{\mathcal{V}\in\Theta^{*}}\mathbb{P}(\mathcal{V}\in\widehat{\Theta}\mid\mathcal{D}_{\text{train}})\geq 1-\alpha.$
As is standard, the guarantee is asymptotic. It is also conditional on the training data, which determines the optimal set $\Theta^{*}$. The condition on the third moment of the loss is mild and typically satisfied in practice.
3.5 MCS Ensembles
Usually, an MCS contains more than one model. In such cases, there is a set of good models whose performance is effectively indistinguishable given the limited amount of observations. In such cases, it is natural to take an ensemble approach. This may have the additional benefit of variance reduction and improved predictive performance, as is well-known from the literature on model averaging (hoeting1999bayesian; burnham2002model). In particular, we propose to define the MCS mixture model
$\widehat{f}_{\widehat{\Theta}}(\bm{z})=\frac{1}{|\widehat{\Theta}|}\sum_{\mathcal{V}\in\widehat{\Theta}}\widehat{f}_{\mathcal{V},\mathcal{D}}(\bm{z}),$
which is straightforward to use in both downstream generative and regression-type tasks.
For regression problems, we adapt the estimating equation approach of Nagler2018 to the MCS mixture model. Specifically, let $\psi_{\beta}$ be a function such that $\mathds{E}[\psi_{\beta}(Y)\mid\bm{X}=\bm{x}]=0$ if and only if $\beta$ is the target of inference. Examples are $\psi_{\beta}(y)=y-\beta$ for the conditional mean $\beta(\bm{x})=\mathds{E}[Y\mid\bm{X}=\bm{x}]$ or $\psi_{\beta}(y)=\mathds{1}\{y<\beta(\bm{x})\}-\tau$ for the conditional $\tau$-quantile. We now approximate the conditional expectation by the fitted model, i.e., we solve
$\int\psi_{\beta}\bigl(y\bigr)\widehat{f}_{\widehat{\Theta}}(y\mid\bm{x})dy=0.$ (2)
In practice, we consider a set of equally spaced grid points $\{y_{1},\dots,y_{G}\}$. If (2) holds, we also have
$\sum_{g=1}^{G}\psi_{\beta}(y_{g})w_{\widehat{\Theta}}(\bm{x},y_{g})\approx 0,$
with
$w_{\widehat{\Theta}}(\bm{x},y)=\widehat{f}_{Y}(y)\sum_{\mathcal{V}\in\widehat{\Theta}}\widehat{c}_{\mathcal{V},\mathcal{D}}(\widehat{F}_{Y}(\bm{x}),\widehat{F}_{Y}(y)),$
a weight that is the product of the copula density and the marginal density of $Y$. This approach is convenient because it puts no restrictions on the structure of the vine copula, facilitates model averaging, and can be used for all sorts of conditional quantities. We refer to Section˜4.1 and Appendix˜C for additional discussions and computational details.
4 Experiments
Table 1: Average NLL (standard error) on test data for density estimation tasks. Best results in bold.
Energy Concrete Airfoil Wine Ccpp California
Dissmann 1.95 (0.35) 7.14 (0.11) 3.25 (0.04) 8.49 (0.19) 6.8 (0.01) 3.63 (0.16)
Kraus 1.51 (0.32) 7.16 (0.12) 3.27 (0.04) 8.39 (0.15) 6.8 (0.01) 3.6 (0.17)
RS-B (50) 0.47 (0.36) 7.08 (0.11) 3.19 (0.06) 8.51 (0.19) 6.77 (0.01) 3.49 (0.17)
RS-B (100) 0.36 (0.36) 7.05 (0.14) 3.20 (0.06) 8.51 (0.19) 6.76 (0.01) 3.47 (0.16)
RS-B (500) 0.20 (0.37) 7.04 (0.11) 3.09 (0.03) 8.51 (0.19) 6.75 (0.01) 3.45 (0.17)
RS-E (50) 0.12 (0.43) 6.55 (0.10) 3.05 (0.05) 8.11 (0.18) 6.74 (0.01) 3.41 (0.16)
RS-E (100) 0.13 (0.34) 6.51 (0.09) 3.03 (0.04) 8.06 (0.19) 6.74 (0.01) 3.36 (0.16)
RS-E (500) -0.28 (0.32) 6.49 (0.09) 3.00 (0.04) 7.92 (0.18) 6.73 (0.01) 3.31 (0.17)
To validate the new methods, we provide a comprehensive comparison using popular real-world data sets on density estimation and regression tasks.
4.1 Setup
Data sets
We use five data sets from the UCI repository (Dua:2019), Energy Efficiency ($n=768$, $p=8$), Concrete Compressive Strength ($n=1030$, $p=8$), Airfoil Self-Noise ($n=1503$, $p=5$), Wine quality ($n=4898$, $p=12$), Combined Cycle Power Plant (Ccpp) ($n=9568$, $p=5$), as well as the California Housing dataset ($n=20640$, $p=8$) (Pace1997). All have a continuous label variables and are commonly used for benchmarking density estimation and tabular regression models. The number of features is in the typical range for applications of vine copula models, see also the discussions in Section˜2.4 and Section˜5 for additional challenges in higher-dimensional settings. For all datasets, we standardize the features to have mean zero and unit variance, randomly split the data into 80% training, and 20% test sets. For methods requiring a validation set, we further split the training set into 75% training and 25% validation.
Computational details
All experiments were run on a standard desktop machine. A Python package implementing the proposed methods and scripts to reproduce the experiments are available in the supplementary materials and described in Appendix˜C. It relies mostly on the pyvinecopulib package (pyvinecopulib) for vine copula modeling. Marginal densities and distribution are estimated by kernel local-likelihood. As for the pair-copulas, we use the nonparametric TLL estimator, which combines state-of-the-art performance with computational efficiency (nagler2017nonparametric).
For MCS inference, we rely on a Python translation of the R (R2025) script gracefully provided by the authors of kim2025locally, improved to scale as $O(Mn)$ instead of $O(M^{2}n)$; see Section˜A.2 for details. All experiments were repeated 10 times with different random seeds, and we report the average performance and standard error across these replications.
4.2 Density Estimation
In this experiment, we assess the quality of density estimation by stacking the features and the label, so that the problem becomes $(p+1)$-dimensional. We compare the following methods:
  • Dissmann: A first benchmark method using the heuristic of Dissmann2013, based on absolute Kendall’s $\tau$.
  • Kraus: A second benchmark method using the heuristic of Kraus2017, based on a combination of Kendall’s $\tau$ and $p$-values of tests for the simplifying assumption.
  • RS-B (M): Our random search Algorithm˜1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(\bm{Z})$.
  • RS-E (M): An ensemble over the MCS obtained by Algorithm˜2 with $M$ candidates with the same loss as RS-B and $\alpha=0.05$ (see Section˜3.5).
Other approaches mentioned in Section˜2.4 are not included in the comparison because of their prohibitive computational cost or lack of general applicability.
In Table˜1, we show the average negative log-likelihood (NLL) on test data. Except for the wine data set, both of our random search approaches outperform the benchmarks, irrespective of the number of candidates $M$. The performance improves with $M$, as expected, and our RS-E (500) method is the best on all data sets, albeit the margin is small in in some cases. The results also explain why the data set Energy is not included in Figure˜1: the performance gain over the benchmark is simply too large to be displayed on the same scale.
Figure 3: Density estimation: Lines indicated how often the Dissmann method is part of the $95\%$-MCS (computed on the training set) across different datasets, estimated over 20 seeds.
Table 2: Average RMSE (standard error) on test data for mean regression tasks. Best results in bold.
Energy Concrete Airfoil Wine Ccpp California
Dissmann 2.76 (0.1) 7.12 (0.13) 4.52 (0.08) 0.63 (0.01) 4.09 (0.03) 0.66 (0.0)
Kraus 3.4 (0.13) 7.09 (0.14) 4.48 (0.1) 0.63 (0.01) 4.26 (0.02) 0.65 (0.0)
RS-B (50) 2.55 (0.23) 6.98 (0.13) 3.73 (0.06) 0.64 (0.01) 4.08 (0.03) 0.62 (0.0)
RS-B (100) 2.33 (0.2) 7.02 (0.15) 3.76 (0.07) 0.64 (0.01) 4.07 (0.03) 0.62 (0.0)
RS-B (500) 2.06 (0.13) 6.9 (0.16) 3.73 (0.07) 0.63 (0.01) 4.08 (0.03) 0.62 (0.0)
RS-E (50) 2.0 (0.06) 6.28 (0.1) 3.73 (0.06) 0.61 (0.01) 4.07 (0.03) 0.6 (0.0)
RS-E (100) 2.04 (0.07) 6.24 (0.1) 3.78 (0.06) 0.61 (0.0) 4.06 (0.03) 0.6 (0.0)
RS-E (500) 1.87 (0.06) 6.24 (0.08) 3.79 (0.07) 0.61 (0.01) 4.06 (0.03) 0.6 (0.0)
A key benefit of the MCS approach is that it enables assessing whether a benchmark is statistically indistinguishable from the best candidates using only the training data. Figure˜3 reports the frequency with which Dissmann belongs to the $95\%$-MCS across datasets, which closely mirrors the results in Table˜1. Notably, for the only dataset where Dissmann is better than RS-B (Wine), it is included in the MCS in every replication. In practice, this provides a simple check: the benchmark can be retained when it is competitive, without risking performance loss when it is clearly inferior.
4.3 Mean and Median Regression
In this experiment, we compare the following methods:
  • Dissmann: The same heuristic as above, albeit combined with Nagler2018 to turn the joint density estimate into a conditional mean or median regression function (see Section˜3.5).
  • Kraus: The method of kraus2017d, which uses the fact that D-Vines admit closed-form conditional distributions.
  • RS-B (M): Our random search Algorithm˜1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(Y|\bm{X})$, turned into a regression function as for Dissmann.
  • RS-E (M): Same, albeit for Algorithm˜2 with the MCS using $\alpha=0.05$, and Nagler2018 adapted to ensembles as described in Section˜3.5.
In Table˜2, we show the root mean squared error (RMSE) on test data for mean regression. Similarly to the density estimation experiment, random search methods outperform the benchmarks in most cases, with performance improving with $M$. Again, RS-E (500) is the best method in five out of six datasets, with the RMSE of RS-E (M) being consistently lower than that of RS-B (M) for almost all $M$ and datasets. The exact same pattern is observed for median regression when evaluated via mean absolute error (MAE); see Appendix˜D for details. This suggests that ensembling via the MCS is particularly beneficial in regression-type problems, where the model selection is not directly optimized for the evaluation metric.
4.4 Probabilistic Forecasting
Table 3: Average CRPS (standard error) on test data for probabilistic forecasting tasks. Best results in bold.
Energy Concrete Airfoil Wine Ccpp California
Dissmann 1.36 (0.03) 3.91 (0.07) 2.38 (0.04) 0.31 (0.0) 2.25 (0.01) 0.33 (0.0)
Kraus 1.41 (0.05) 3.94 (0.06) 2.37 (0.04) 0.31 (0.0) 2.38 (0.01) 0.33 (0.0)
RS-B (50) 1.14 (0.07) 3.81 (0.08) 1.95 (0.02) 0.31 (0.0) 2.24 (0.01) 0.31 (0.0)
RS-B (100) 1.12 (0.1) 3.8 (0.07) 1.97 (0.03) 0.31 (0.0) 2.23 (0.01) 0.31 (0.0)
RS-B (500) 0.97 (0.05) 3.78 (0.08) 1.94 (0.03) 0.3 (0.0) 2.24 (0.01) 0.3 (0.0)
RS-E (50) 0.93 (0.02) 3.39 (0.05) 1.94 (0.03) 0.29 (0.0) 2.24 (0.01) 0.29 (0.0)
RS-E (100) 0.95 (0.02) 3.37 (0.05) 1.97 (0.03) 0.29 (0.0) 2.24 (0.01) 0.29 (0.0)
RS-E (500) 0.89 (0.02) 3.37 (0.04) 1.97 (0.03) 0.29 (0.0) 2.23 (0.01) 0.29 (0.0)
In this experiment, we compare the same methods as in the regression experiment. However, we evaluate the quality of the full predictive distribution via the continuous ranked probability score (CRPS) (gneiting2014probabilistic), a popular proper scoring rule for probabilistic forecasting. Specifically, we use the fact that the methods from the previous experiment yield quantiles at all levels, and compute the CRPS via
$\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(y-F^{-1}(\tau))d\tau,$
with $\rho_{\tau}(u)=u(\tau-\mathds{1}\{u<0\})$ the check function. Results are shown in Table˜3. Outperformance of our random search methods over the benchmarks is even more pronounced in this experiment, and ensembles via the MCS are again particularly effective.
4.5 Runtime
In Figure˜4, we show the relative CPU time of our random search methods compared to Dissmann on the Concrete data set. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time grows with $M$, since some fraction of the candidates are retained in the MCS and used for prediction.
While the computational cost of the random search methods is higher than that of Dissmann, the total cost is still tiny for real datasets of this scale. For instance, training and inference average respectively 0.357 and 0.006 seconds on a single CPU core for the Concrete data set. Additionally, the random search methods are easily parallelizable, since candidates can be fitted independently. See Appendix˜D for additional runtime results on other data sets.
Figure 4: Relative CPU time of random search methods compared to Dissmann on the Concrete data set.
5 Conclusion
We propose a simple and yet effective method for learning vine copula structures based on random search and model confidence sets. Our experiments on real-world data sets demonstrate that the new method outperforms existing greedy heuristics in both density estimation and regression tasks, often by a large margin.
A limitation of the method is the additional computational cost compared to greedy heuristics. However, this cost is modest if the dimension is not too high and easily parallelizable, making it a worthwhile trade-off for the significant performance gains achieved. In truly high-dimensional problems, however, any structure learning method must be combined with additional sparsity-inducing mechanisms, such as truncation, regularization, or variable selection. While beyond the scope of the current work, designing more sophisticated sampling methods that incorporate such mechanisms is an important and promising direction for future research.
Acknowledgements.
We thank Johanna F. Ziegel for pointing us to the literature on model confidence sets. This work was partially supported by the Hasler Stiftung under grant number 2025-05-01-519.
Appendix A DETAILS ON THE MCS ALGORITHM
In the following, we recall the relevant details of the MCS algorithms given by kim2025locally to keep our paper self-contained. Recall the notation of our main paper, specifically Algorithm˜2: There are $M$ candidate models $\{\mathcal{V}_{m}\}_{m=1}^{M}$ and we have already computed validation losses $L_{i,m}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$. For ease of notation, define the loss differences $\Delta_{m,k}^{(i)}=L_{i,m}-L_{i,k}$ for $1\leq i\leq n_{\text{val}}$ and $1\leq m,k\leq M$.
A.1 The DA Test Statistic and DA-MCS-marg Construction
We first describe the discrete argmin (DA) hypothesis test introduced by kim2025locally which serves as the foundation for constructing our MCS. The basic algorithm to compute the test statistic from a set of loss observations is given in Algorithm˜A.3.
Algorithm A.3 DA test statistic
1:Loss observations $\{L_{i,m}\colon 1\leq i\leq N,\,1\leq m\leq M\}$, target index $r\in\{1,\dots,M\}$.
2:Test statistic $\widehat{T}_{r}$.
3:Define a split of indices: $I_{1}=\{1,\dots,\lfloor N/2\rfloor\}$ and $I_{2}=\{\lfloor N/2\rfloor+1,\dots,N\}$.
4:On $I_{1}$, find the alternative model that minimizes the cumulative validation loss:
$\widehat{m}_{-r}\in\operatorname*{arg\,min}_{m\in\{1,\dots,M\}\setminus\{r\}}\sum_{i\in I_{1}}L_{i,m},$
breaking ties by the smallest index.
5:On $I_{2}$, compute
$\widehat{T}_{r}\;=\;\frac{1}{\widehat{\sigma}_{r,\widehat{m}_{-r}}}\cdot\frac{1}{\sqrt{|I_{2}|}}\sum_{i\in I_{2}}\Delta_{r,\widehat{m}_{-r}}^{(i)},$
where $\widehat{\sigma}_{r,\widehat{m}_{-r}}^{2}$ is the sample variance of $\{\Delta_{r,\widehat{m}_{-r}}^{(i)}:i\in I_{2}\}$.
Under the null hypothesis that model $r$ has the smallest expected out-of-sample loss among the $M$ candidates, a one-sided test rejects for large positive values of $\widehat{T}_{r}$, e.g. when $\widehat{T}_{r}>\Phi^{-1}(1-\alpha)$, where $\Phi$ is the standard normal CDF.
To obtain an MCS with marginal coverage guarantees, one can simply invert the DA test: compute $\widehat{T}_{r}$ for all $r=1,\dots,M$ by running Algorithm˜A.3 and define
$\widehat{\Theta}=\left\{r\in\{1,\dots,M\}\colon\;\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\right\}.$
This procedure, which we call DA-MCS-marg, yields an MCS with marginal coverage $1-\alpha$ under mild conditions, as formally stated in Proposition˜3.2.
A.2 Efficient Implementation and Complexity
When implemented exactly as described above, the DA-MCS-marg procedure has complexity $O(n_{\text{val}}M^{2})$:
  • Each call to Algorithm˜A.3 requires $O(n_{\text{val}}M)$ time to compute the cumulative losses, $O(M)$ time to find the argmin, and $O(n_{\text{val}})$ time to compute the sample variance and test statistic. This results in a total of $O(n_{\text{val}}M)$ time per call.
  • Because Algorithm˜A.3 is called $M$ times in total for DA-MCS-marg, the overall complexity is $O(n_{\text{val}}M^{2})$
However, this complexity can be reduced to $O(n_{\text{val}}M)$ for both procedures by avoiding redundant computation when looping over $r\in\{1,\dots,M\}$:
  • The column sums $\overline{L}_{m}=\sum_{i\in I_{1}}L_{i,m}$ for $1\leq m\leq M$ can be computed once in $O(n_{\text{val}}M)$ time and accessed in $O(1)$ time in each call to Algorithm˜A.3.
  • The argmins $\widehat{m}_{-1},\dots,\widehat{m}_{-M}$, can be computed simultaneously in $O(M)$ time. Specifically, one pass through the precomputed column sums $\overline{L}_{1},\dots,\overline{L}_{M}$ suffices to find the two smallest numbers $\overline{L}_{j_{1}}$ and $\overline{L}_{j_{2}}$ among $\{\overline{L}_{1},\dots,\overline{L}_{M}\}$ (breaking ties by index) along with corresponding indices $j_{1}$ and $j_{2}$. Then, use another pass to set $\widehat{m}_{-r}=j_{1}$ if $\overline{L}_{r}\neq\overline{L}_{j_{1}}$ and $\widehat{m}_{-r}=j_{2}$ otherwise, for all $r=1,\dots,M$.
Figure˜5 compares the average computation time of the naive $O(n_{\text{val}}M^{2})$ implementation and the fast $O(n_{\text{val}}M)$ implementation of DA-MCS-marg (as well as a variant for obtaining uniform MCS) on a log-log scale.
Figure 5: Average computation time of different DA-MCS Python implementations.
A.3 Runtime Comparison of MCS Methods
In Figure˜6, we compare the average computation time of different MCS methods. Method names are from the R script provided by the authors of kim2025locally; KR-PLG refers to DA-MCS-marg. We observe that the method is significantly faster than competing methods.
Figure 6: Average computation time of MCS methods.
Appendix B PROOFS OF THEORETICAL RESULTS
The proof uses similar arguments as kim2025locally, but requires some technical modifications to adapt to our setting. We start with a small lemma simplifying the moment condition required by kim2025locally. To simplify notation, we write $\mathds{E}^{\prime}[\cdot]=\mathds{E}[\cdot\mid\mathcal{D}_{\text{train}}]$ and $\mathbb{P}^{\prime}(\cdot)=\mathbb{P}(\cdot\mid\mathcal{D}_{\text{train}})$ throughout this section.
Lemma B.1.
Let $M\in\mathds{N}$ and define $\Delta_{m,k}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})-L(\widehat{f}_{\mathcal{V}_{k},\mathcal{D}_{\text{train}}},\bm{Z})$ and $\overline{\Delta}_{m,k}=\Delta_{m,k}-\mathds{E}^{\prime}[\Delta_{m,k}]$. If
$\limsup_{n\to\infty}\max_{1\leq m\leq M}\mathds{E}^{\prime}[|L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}]<\infty$ (3)
there is $K\in(0,\infty)$ such that for all $n$ large enough,
$\max_{1\leq m,k\leq M}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}\leq K,$ (4)
with the convention $0/0=1$.
Proof.
The moment condition (3) implies that there are $C\in(0,\infty)$ and $n_{0}\in\mathds{N}$ such that for all $n\geq n_{0}$,
$\max_{1\leq m\leq M}\mathds{E}^{\prime}[|L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}]\leq C.$
Suppose $n\geq n_{0}$ from now on. It holds
$\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]$ $\leq\mathds{E}^{\prime}[(|\Delta_{m,k}|+|\mathds{E}^{\prime}[\Delta_{m,k}]|)^{3}]\leq 4\mathds{E}^{\prime}[|\Delta_{m,k}|^{3}]+4|\mathds{E}^{\prime}[\Delta_{m,k}]|^{3}\leq 8\mathds{E}^{\prime}[|\Delta_{m,k}|^{3}],$
using triangle inequality in the first, $(a+b)^{3}\leq 4(|a|^{3}+|b|^{3})$ in the second, and Jensen’s inequality in the last step. Using the same arguments, we also have
$\mathds{E}^{\prime}[|\Delta_{m,k}|^{3}]$ $\leq 4\mathds{E}^{\prime}[|L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}]+4\mathds{E}^{\prime}[|L(\widehat{f}_{\mathcal{V}_{k},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}]\leq 8C.$
Overall, we have shown that $\max_{1\leq m,k\leq M}\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]\leq 64C$. Now let
$\mathcal{S}=\left\{(m,k)\in\{1,\dots,M\}^{2}\colon E^{\prime}[|\overline{\Delta}_{m,k}|^{2}]>0\right\},$
$\mathcal{S}^{c}=\{1,\dots,M\}^{2}\setminus\mathcal{S}$, and define $c=\min_{(m,k)\in\mathcal{S}}E^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}$, noting that $c>0$. We have
$\max_{1\leq m,k\leq M}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}$ $\leq\max_{(m,k)\in\mathcal{S}}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}+\max_{(m,k)\in\mathcal{S}^{c}}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}$
$\leq 64C/c+\max_{(m,k)\in\mathcal{S}^{c}}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}.$
Recall that $(m,k)\in\mathcal{S}^{c}$ implies $\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]=0$. This further implies $\overline{\Delta}_{m,k}=0$ almost surely, and because $\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]<\infty$, it follows that $\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]=0$. Altogether have shown that
$\max_{1\leq m,k\leq M}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}}\leq 64C/c+1=:K<\infty,$
which completes the proof. ∎
Proof of Proposition˜3.2.
Let $r\in\Theta^{*}$ be arbitrary but fixed. Define
$\widehat{T}_{r}^{*}=\frac{1}{\widehat{\sigma}_{r,\widehat{m}_{-r}}}\cdot\frac{1}{\sqrt{|I_{2}|}}\sum_{i\in I_{2}}\left(\Delta_{r,\widehat{m}_{-r}}^{(i)}-\mathds{E}^{\prime}[\Delta_{r,\widehat{m}_{-r}}^{(i)}\mid\widehat{m}_{-r}]\right),$
a centered version of $\widehat{T}_{r}$ that satisfies $\widehat{T}_{r}\leq\widehat{T}_{r}^{*}$ since $r\in\Theta^{*}$ implies $\mathds{E}^{\prime}[\Delta_{r,\widehat{m}_{-r}}^{(i)}\mid\widehat{m}_{-r}]\leq 0$. It holds
$\mathbb{P}^{\prime}(r\in\widehat{\Theta}\mid\widehat{m}_{-r})$ $=\mathbb{P}^{\prime}(\widehat{T}_{r}\leq\Phi^{-1}(1-\alpha)\mid\widehat{m}_{-r})$
$\geq\mathbb{P}^{\prime}(\widehat{T}_{r}^{*}\leq\Phi^{-1}(1-\alpha)\mid\widehat{m}_{-r})$
$\geq 1-\alpha-\sup_{t\in\mathds{R}}\left|\mathbb{P}^{\prime}(\widehat{T}_{r}^{*}\leq t\mid\widehat{m}_{-r})-\Phi(t)\right|.$
To show that the supremum on the right is negligible, we apply the Berry-Esseen bound for Student’s statistic of bentkus1996berry, conditionally on $\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M}$, and $\widehat{m}_{-r}$ (which are all independent of $\{L_{i,m}\colon n_{\text{val}}/2<i\leq n_{\text{val}}\}$). This yields
$\sup_{t\in\mathds{R}}\left|\mathbb{P}^{\prime}(\widehat{T}_{r}^{*}\leq t\mid\widehat{m}_{-r})-\Phi(t)\right|\leq\frac{C}{\sqrt{n_{\text{val}}}}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{r,\widehat{m}_{-r}}|^{3}\mid\widehat{m}_{-r}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{r,\widehat{m}_{-r}}|^{2}\mid\widehat{m}_{-r}]^{3/2}}\leq\frac{C}{\sqrt{\lfloor\eta n\rfloor}}\max_{1\leq m,k\leq M}\frac{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{3}]}{\mathds{E}^{\prime}[|\overline{\Delta}_{m,k}|^{2}]^{3/2}},$
for a universal constant $C\in(0,\infty)$. Lemma˜B.1 implies that the right-hand side is bounded by $CK/\sqrt{\lfloor\eta n\rfloor}$ for all $n$ large enough. Altogether, this gives
$\mathbb{P}^{\prime}(r\in\widehat{\Theta})=\mathds{E}^{\prime}[\mathbb{P}^{\prime}(r\in\widehat{\Theta}\mid\widehat{m}_{-r})]\geq 1-\alpha-\frac{CK}{\sqrt{\lfloor\eta n\rfloor}}=1-\alpha-o(1),$
as claimed. ∎
Appendix C IMPLEMENTATION DETAILS
To ensure both reproducibility and impact, we bundle the proposed methods into a Python package named vinesforests, whose estimators follow the scikit-learn (scikit-learn) API (fit, predict, score, score_samples, etc.) and are compatible with numpy (numpy) arrays and, where indicated, pandas (pandas) DataFrames. As part of the supplementary material, we provide a zip file containing the following file structure:
  • README.md — Instructions for installation and running experiments.
  • pyproject.toml — Configuration file for the package, Pixi and dependencies.
  • src/ — Source code for the vinesforests package.
    • vine: This module exposes two estimator classes, VineDensity and VineRegressor, which implement single vine density estimation and regression, respectively.
    • forest: This module exposes two estimator classes, VineForestDensity and VineForestRegressor, which implement the ensemble methods for density estimation and regression, respectively.
    • benchmark: This module exposes two estimator classes, KrausVineDensity and KrausVineRegressor, which are wrappers around R implementations of the baselines.
    • experiments: This module contains utilities for running experiments, including a CLI interface.
    • mcs.py — Implementation of the MCS procedures.
  • tests/ — Unit tests.
  • notebooks/ — Scripts and Jupyter notebooks allowing to fully reproduce our experiments and analysis.
C.1 Single vine estimators
The class VineBase, which derives from sklearn.base.BaseEstimator, centralizes preprocessing and vine–copula fitting using pyvinecopulib (pyvinecopulib). Inputs may be numpy arrays or pandas DataFrames containing numeric columns, ordered categoricals, and unordered categoricals. Unordered categoricals are expanded to dummy variables via a map $x\mapsto\{\mathbf{1}(x=\ell_{j})\}_{j>1}$, with dummies represented as ordered categoricals taking values $\{0,1\}$. Note that, for both VineBase and its subclasses VineDensity and VineRegressor, expensive computations are batch vectorized and processed in chunks of size batch_size to trade memory for throughput.
Marginal distributions
A per-feature univariate kernel density estimator, pyvinecopulib.Kde1d, is used for each marginal; its type is set to "continuous" or "discrete" according to the post-expansion schema. We refer to this notebook with examples and the pyvinecopulib.Kde1d documentation for more details, notably on automatic bandwidth selection. For a given model, the pseudo–observations used by the vine are computed as follows. For continuous variables, the probability integral transform $U=F(Z)$ given by pyvinecopulib.Kde1d.cdf is used. For discrete features, both $F(Z)$ and the left–limit $F(Z^{-})$ are computed and stacked.
Vine copula
Given marginals $\{F_{j}\}_{j}$, the vine copula $C$ is fit on pseudo–observations $U=(U_{1},\ldots,U_{d})$ with $U_{j}=F_{j}(Z_{j})$. In the density context, $\bm{Z}=\bm{X}$ is the feature vector. In the regression context, $\bm{Z}=(Y,\bm{X})$ is the joint vector of target and features, with $Y$ in the first dimension treated as continuous. For vine copulas, we use the pyvinecopulib.Vinecop class and refer to the documentation for details, as well as the notebook with examples. In vinesforest, a vine copula is estimated via pyvinecopulib.Vinecop.from_data(data=U, structure=..., var_types=..., controls=...):
  • If a structure is not specified, pyvinecopulib performs structure selection using Dissmann2013 by default. Otherwise, it needs to be an instance of pyvinecopulib.RVineStructure (e.g., generated via pyvinecopulib.RVineStructure.simulate(d)).
  • By default, the controls are pyvinecopulib.FitControlsVinecop(family_set=[pyvinecopulib.tll], num_threads=1), so that the TLL family is used for all pair-copulas. Other families can be specified via the family_set argument, parallelism is controlled with num_threads, and we refer to the documentation for more options.
This class provides a _pdf_samples(X, y=None, log=False, copula_only=False) method which is reused by subclasses. It implements the pipeline of computing pseudo–observations, evaluating the copula density via pyvinecopulib.Vinecop.pdf, and potentially multiplying by marginal densities using pyvinecopulib.Kde1d.pdf.
VineDensity
For this class, which derives from VineBase, the fit method calls the above pipeline on $X$ only. The log-likelihood per sample is returned by score_samples(X), and the average by score(X). Sampling draws $U\sim C$ via vine.simulate and transforms back with the fitted marginals using $X_{j}=F_{j}^{-1}(U_{j})$.
VineRegressor
For this class, which derives from VineBase as well, the fit fits the vine on the joint $(Y,X)$. Predictions via predict use conditional importance weights derived from the copula density. For a test observation $x$, the method computes
$w_{i}(x)\propto\begin{cases}c_{Y,X}(U_{Y}(y_{i}),U_{X}(x)),\quad i=1,\dots,n_{\text{train}},&\text{if }\texttt{use\_grid = False},\\ c_{Y,X}(U_{Y}(y_{i}),U_{X}(x))f_{Y}(y_{i}),\quad i=1,\dots,n_{\text{grid}},&\text{if }\texttt{use\_grid = True},\end{cases}$
where the grid is a set of equally spaced grid points $\{y_{1},\dots,y_{G}\}$ in $Y$-space, and $U_{Y}(y_{i})=F_{Y}(y_{i})$ is computed with the marginal CDF. The second option is usually faster when the training set is large, as it reweights by the marginal density $f_{Y}(y)$ to correct for non-uniformity of grid points instead of using the entire empirical distribution. Conditional expectations are computed in closed form via $\sum_{i}w_{i}(x)y_{i}$, and conditional quantiles are computed by weighted quantiles of $\{y_{i}\}_{i}$ using the “inverted CDF” definition. For more details, we refer to the numpy.quantile definition, and in particular its weights argument.
To evaluate the conditional log-likelihood used in the forest selection (see below), VineRegressor provides a method copula_marginal_density(X, log=False, n_grid=101), which uses the identity
$\log f_{Y\mid X}(y\mid x)=\log c_{Y,X}(U_{Y}(y),U_{X}(x))-\log c_{X}(U_{X}(x))+\log f_{Y}(y),$
where $c_{X}(u_{X})=\int_{0}^{1}c_{Y,X}(u_{Y},u_{X})\,du_{Y}$ is approximated with Simpson’s rule on a uniform grid over $[0,1]$.
C.2 Forest estimators (ensemble level)
Similarly as in the single vine case, a class VineForestBase deriving from sklearn.base.BaseEstimator implements structure randomization, bootstrap resampling, validation splitting, survivor (either via the MCS or simply those better than Dissmann2013 in the validation set) selection, and parallel fitting and prediction averaging. It uses a base learner, which can be either VineDensity or VineRegressor, and has two subclasses VineForestDensity and VineForestRegressor. After fitting the default estimator using (Dissmann2013) for structure selection, for a requested $M$ base learners, VineForestBase proceeds as follows:
  1. 1.
    For each $r=1,\ldots,M$, use pyvinecopulib.RVineStructure.simulate to simulate a random $R$-vine structure $\mathcal{V}_{r}$, optionally bootstraps the training rows, and fits a cloned base estimator with this structure.
  2. 2.
    Scores each candidate on the held-out validation fold using per-sample log-likelihoods: $\log f_{X}(x)$ for density and $\log f_{Y\mid X}(y\mid x)$ for regression (as above).
  3. 3.
    Selects survivors either by keeping all candidates strictly improving over the default or by running the MCS procedure at level $\alpha$ (da_mcs_marg). Optionally best_only retains the single best survivor.
  4. 4.
    Refits all survivors on the full training set and averages predictions at inference time (log of mean densities for VineForestDensity; averaging conditional weights for VineForestRegressor).
Note that:
  • Parallelism uses joblib (joblib). When n_jobs exceeds the number of survivors, the code increases the pyvinecopulib num_threads of each base learner to utilize all cores.
  • The selection routine da_mcs_marg (module mcs.py) implement the procedures described in Appendix˜A, with the refinement described in Section˜A.2 to make it scale linearly with the number of candidates.
  • The two derived classes VineForestDensity and VineForestRegressor only implement respectively the score/score_sample, and predict methods, which call the appropriate methods of the base learners and aggregate results.
C.3 Other Components
Determinism and Performance
All randomization is seeded through a user–provided integer seed; random vine structures derive from numpy RNG streams. Batch size controls the memory/runtime trade-off for likelihood evaluation and prediction. Vectorization is used throughout, and heavy loops are limited to batched calls into pyvinecopulib. Where feasible, computations are parallelized via joblib and intra-estimator threads in pyvinecopulib.
Benchmarks and R Interfacing
The classes KrausVineRegressor and KrausVineDensity are wrappers for the R packages vinereg and kdevine, respectively, through rpy2. A small mixin, RInterface, ensures silent R output, proper conversion of pandas categoricals to R factors, and robust conversion between numpy/pandas and R objects. The two classes are also derived from VineRegressor and VineDensity, respectively, to ensure compatibility with the scikit-learn API and the rest of the codebase.
Experiments
The file utils.py offers: (i) expand_factors for categorical handling; (ii) crps_from_quantiles implementing Simpson–rule CRPS from predictive quantiles. As for the file real_data.py, it contains lightweight loaders for the UCI/OpenML datasets used in the paper and applies standardization where appropriate; when categorical variables are present (e.g., Bike Sharing), ordered categoricals are preserved for compatibility with VineRegressor. The rest of the module contains a CLI interface to run experiments, which is used in the provided Jupyter notebooks to reproduce results.
Appendix D ADDITIONAL EXPERIMENTS
D.1 Results for Median Regression
In Table˜4, we report the average mean absolute error (MAE) on test data for median regressions using the same experimental setup as in our main paper. We observe similar trends as for mean regression and probabilistic forecasting. Our random search methods outperform the greedy Dissmann and Kraus algorithms, with RS-E generally performing best.
Table 4: Average mean absolute error (standard error) on test data for median regressions. Best results in bold.
Energy Concrete Airfoil Wine Ccpp California
Dissmann 1.85 (0.06) 5.38 (0.1) 3.31 (0.05) 0.43 (0.01) 3.15 (0.01) 0.46 (0.0)
Kraus 1.93 (0.08) 5.55 (0.08) 3.3 (0.06) 0.44 (0.01) 3.32 (0.01) 0.45 (0.0)
RS-B (50) 1.56 (0.11) 5.35 (0.12) 2.71 (0.04) 0.43 (0.01) 3.14 (0.02) 0.43 (0.0)
RS-B (100) 1.59 (0.2) 5.3 (0.11) 2.75 (0.05) 0.43 (0.01) 3.13 (0.02) 0.42 (0.0)
RS-B (500) 1.32 (0.1) 5.28 (0.12) 2.69 (0.04) 0.42 (0.01) 3.15 (0.02) 0.42 (0.0)
RS-E (50) 1.15 (0.04) 4.74 (0.09) 2.69 (0.04) 0.4 (0.0) 3.13 (0.02) 0.4 (0.0)
RS-E (100) 1.15 (0.04) 4.71 (0.08) 2.75 (0.05) 0.4 (0.0) 3.13 (0.02) 0.41 (0.0)
RS-E (500) 1.05 (0.02) 4.7 (0.07) 2.75 (0.05) 0.4 (0.0) 3.12 (0.01) 0.4 (0.0)
D.2 Runtime Comparison on Other Datasets
In Figure˜7, we show the relative CPU time of our random search methods compared to Dissmann on all data sets. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time can grow with $M$, since some fraction of the candidates are retained in the MCS and used for prediction. However, this increase is not systematic, as there can sometimes be a single (or a few) model(s) in the MCS, irrespective of $M$.
In Table˜5, we report the median CPU time for training and inference using a single vine on all data sets. Except for the California data set, training and inference using a single vine is very fast, taking at most around one second for training and less than 0.5 seconds for inference on a single CPU core. Note that, for inference, the reported CPU time is for predicting the mean, median, and 101 quantiles (from 0% to 100% with a step of 1%) of the conditional distribution of the response given the features at all test points. With around 20’000 data points, the California data set is the largest one considered in our experiments, and for such sizes we would need to optimize how pair-copulas are fitted to further reduce the computational cost. This can be done, for instance, by using a smaller TLL grid size initially (i.e., before applying the MCS), and then re-fitting the selected candidates with a larger grid size.
Figure 7: Relative CPU time of random search methods compared to Dissmann on all data sets.
Table 5: Median CPU time (in seconds) for training and inference using a single vine.
Energy Concrete Airfoil
Training Inference Training Inference Training Inference
Density 0.311 0.005 0.367 0.006 0.204 0.004
Regression 0.382 0.33 0.415 0.453 0.213 0.313
Wine Ccpp California
Training Inference Training Inference Training Inference
Density 0.644 0.015 0.742 0.01 6.141 0.047
Regression 0.659 1.141 1.011 1.531 5.109 7.995
D.3 Generating vines by local perturbations
In this appendix, we report additional experiments comparing uniform random search to a locally perturbed version of Dissmann’s algorithm. Specifically, instead of selecting the maximum spanning tree (MST) obtained from the empirical Kendall’s $\tau$ matrix at each tree level, we generate random trees according to distribution
$P(T)\propto\prod_{(j,k)\in T}|\tau_{j,k}|,$
with the MST being the mode. Uniform sampling from this distribution can be done efficiently via Wilson’s loop-erased random walk algorithm (wilson1996generating), implementented e.g. in boost (boost) and made available in the vine selection context by pyvinecopulib (pyvinecopulib).
Table 6: Additional results for the experiment in Table˜1 for a variant of Algorithm˜2 that generates vines by sampling local perturbations of the Dissmann structure.
Energy Concrete Airfoil Wine Ccpp California
RS-B (50) 1.54 (0.28) 7.1 (0.17) 3.11 (0.06) 8.47 (0.27) 6.77 (0.01) 3.56 (0.28)
RS-B (100) 1.17 (0.4) 7.11 (0.16) 3.1 (0.07) 8.47 (0.27) 6.77 (0.01) 3.57 (0.28)
RS-B (500) 0.48 (0.67) 6.99 (0.18) 3.11 (0.06) 8.47 (0.27) 6.76 (0.01) 3.58 (0.28)
RS-E (50) 0.53 (0.37) 6.63 (0.13) 3.06 (0.07) 8.19 (0.33) 6.75 (0.01) 3.51 (0.26)
RS-E (100) 0.19 (0.33) 6.61 (0.12) 3.06 (0.07) 8.09 (0.33) 6.75 (0.01) 3.49 (0.27)
RS-E (500) -0.11 (0.36) 6.57 (0.15) 3.07 (0.07) 7.8 (0.21) 6.74 (0.01) 3.46 (0.27)
Table˜6 reports results for the experiment in Table˜1 of the main text. We see that uniform random search slightly outperforms the local perturbation. From a theoretical standpoint, there is no reason to believe that a greedy approach will find a good structure. In fact, it is easy to simulate data where greedy methods (and perturbed variants) fail catastrophically: construct a vine where the first tree consists of weak dependencies, while the later trees contain strong dependencies and asymmetric relationships. The greedy method will pick a structure that severely violates the simplified vine assumption (see Section 2.3), and many randomly sampled structures lead to better performance. However, alternative local methods may be interesting to explore in future work.